OpenStudy (anonymous):

Stacy made a deposit of $8,500.00 into an account. After ten years, the account had earned $5,525.00 in interest. What simple interest rate did the account earn?

OpenStudy (quietus):

5525 / 10 = 552.5 (simple interest earned per year)

OpenStudy (quietus):

Please continue on your own.

OpenStudy (anonymous):

No Plz I Need This Now My Final In 20 Mins

OpenStudy (quietus):

since Stacy earned $552.50 per year, you can take 552.5/8500

OpenStudy (quietus):

and there's your answer

OpenStudy (quietus):

0.065

OpenStudy (quietus):

or 6.5%

OpenStudy (anonymous):

You can use the formula, I = PRT to solve. I is interest, P is principal amount, R is rate, and T is time. Now we need to know the rate and we have everything else. So we do I/PT = R so all you do is 5525/8,500*10 The answer should be 6.5 percent. I hope giving you the explanation is useful.

OpenStudy (quietus):

For compound interest please use this formula : M = P( 1 + i )^n M is the final amount including the principal. P is the principal amount. i is the rate of interest per year. n is the number of years invested.

OpenStudy (anonymous):

Thank you silca